Find the equation of the line tangent to the circle x2 + y2 = 25 at point (3,4). Use the general equation of the line for your final answer.
Answer by Boreal(15235)

implicit differentiation relative to x
2x+2y dy/dx=0
2y dy/dx=-2x
dy/dx=-x/y
The slope of the line is -3/4
the point is (3, 4)
y-y1=m(x-x1), m slope (x1, y1) point
y-4=(-3/4)(x-3)
y=-(3x/4)+(25/4)